OpenStudy (anonymous):

5x + 8y = 40 in slope intercept form

OpenStudy (texaschic101):

slope intercept is y = mx + b 5x + 8y = 40 --- subtract 5x from both sides 8y = -5x + 40 -- divide both sides by 8 y = -5/8x + 40/8 y = -5/8x + 5 <== slope intercept form